A rectangular sheet of metal having dimensions 20 cm by 12 cm has squares removed from each of the four corners and the sides bent upwards to form an open box. Determine the maximum possible volume of the box.
Answer #1

Que! dimensions 2o cm by 12 cm A rectangulas having by sheet of metal dimensions 20 cm: 12cm has Squares removed from lach Co스 V = 4123-1672 +604). Maximize the expression Now So for the maxima minima find V(x) ? V(x) = 4( 37?- 32* +60] find V(1).V(a) 16 + 2/19 + [6* 32 3 +32] -32] 4 [ 4 16 - 20 3 -32 4 [3? 32 + 4v13 -32 -3?] 4 ( 32 411g -32] 4 X 4 Tg 16ģ 11 - 4x (-411Vio 4.3588 4.369 4-36 51 2STÁ 2x 4.36 8.72 - 16- 8.72 7.28 li 2 V = 4[ 920 (2.42613 16(2.42612 +60 (2.426) 6)] 4 (14.278 -16